Grade upto college level Electric Current

A block of weight W produces an extension of 9cm when it is hung by an elastic spring of length 60cm and is in equlbrm.The spring is cut into 2 parts , one of length 40cm and the other 20cm they are connected in parallel.The same load W hangs in equlbrmsupported by both. The extension now is

a) 9cm b)6cm c) 3cm d)2cm

plzz give the full method rates assured

Kevin Nash
12 Years agoGrade upto college level

1 Answer

Manoj
8 Years ago
Spring constant for original spring W=k×9K is inversely proportional to length so spring constant for spring is 40cm =3/2kAnd fr 20cm length is 3k3k/2 + 3k =9k/2W=9/2kx9k=9/2kxX=2